OpenStudy (anonymous):

Use the rules for integer addition to simplify the expression. -5 + (+7) =

OpenStudy (tiffanymak1996):

-5+7=7-5 can u do the rest?

ganeshie8 (ganeshie8):

a minus sign before 5 : -5

ganeshie8 (ganeshie8):

a plus sign before 7 : +7

OpenStudy (anonymous):

yes does that mean its negative

ganeshie8 (ganeshie8):

yes !

OpenStudy (tiffanymak1996):

yes, but the final answer's not

OpenStudy (anonymous):

so the answer would be +2

ganeshie8 (ganeshie8):

if the signs are used before, they represent whether the quantity is positive or negative. it got nothing to do wid subtraction or addition

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

-5 : negative 5 +5 : positive 5

OpenStudy (anonymous):

so when its -5 + (+7) = you would just go -5 + 7 and you get your answer?

ganeshie8 (ganeshie8):

subtraction/addition are binary operations. they require two numbers. when the - or + symbol is used only on a number, they represent the number

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

so why is the ( and ) there are they to just let you know that it is a positive number?

ganeshie8 (ganeshie8):

-5 + (+7) = (-5) + (+7) = -5 + 7

ganeshie8 (ganeshie8):

yes paranthesis is just for maintaining clarity here

ganeshie8 (ganeshie8):

see the second line, how neatly it says, you need to add two numbers

OpenStudy (anonymous):

ok so for -7 + (-4) =

OpenStudy (anonymous):

would it be -11

ganeshie8 (ganeshie8):

yes !!

OpenStudy (anonymous):

o ok this is easy

ganeshie8 (ganeshie8):

-7 + (-4) = (-7) + (-4) = -7 - 4

ganeshie8 (ganeshie8):

plus * minus = minus

ganeshie8 (ganeshie8):

we used that when removing paranthesis

OpenStudy (anonymous):

ok wait i need help with one more

OpenStudy (anonymous):

+12 + (-11) =

OpenStudy (anonymous):

would it be +1

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

let me knw if you need a practical analogy for this. il try and provide....

ganeshie8 (ganeshie8):

like, (+12) + (+1) means, adding 1 apple to 12 apples

ganeshie8 (ganeshie8):

(+12) + (-1) what does it mean ?

OpenStudy (anonymous):

ok i think i know this one just want to be 100% ok

OpenStudy (anonymous):

Explain how to find the sum -64 + (+41) using a number line, then give the answer.

OpenStudy (anonymous):

-65 + 41 = -24

ganeshie8 (ganeshie8):

wait dont jump

ganeshie8 (ganeshie8):

it asks for answer, thats right!

OpenStudy (anonymous):

so is that right?

ganeshie8 (ganeshie8):

but it also asks to work out the answer on number line

ganeshie8 (ganeshie8):

i think you got the answer quickly by doing this : 1) i have to do -65+41 2) let me first find out, 65-41 3) then il add - sign to the answer 2a) 65-41 = 24 3a) so, the answer is -24

ganeshie8 (ganeshie8):

right ?